# Bootstrapping non-unitary CFTs

### v1

@@ III.1 Search for exact spectra with $c<1$ CFTs
Table 2: Search window for 33-state search
@@
III.2 Result for search in truncated space with $c<1$ CFTs
In the following, we utilize our algorithms in a region where 5 internal state processes can be found. The search window of this region is given in Table 3. We start with a search of one state and keep adding more states until the search finds points above the first state reward threshold. The maximum first state reward for different $N$ is given in Table 4, where we see that for $N<4$ one cannot surpass the threshold of 10. Starting at $N=4$, the threshold is passed and the points found are close to the first four states of known processes. We present five examples in Table 5 Appendix B. This demonstrates that our method can be used to find approximate truncated solutions. If we attempt a six state search, we find no improvement in reward compared to the exact search as indicated in Table 4. Furthermore, the mean OPE coefficient of the last state is of the same order of magnitude as its standard deviation (both around $\sim 10^{-8}$), therefore undistinguishable from zero. The other parameters for the best five and six state solutions are within $4\%$ in relative terms, as shown in Table 6. Thus we conclude that the six state search really finds a five state with the extra state having an OPE of near zero.
Table 3: Search window for 55-state search
@@ III.2 Result for search in truncated space with $c<1$ CFTs
$h_{int,1}$ $\left[0.23,0.3\right]$
$h_{int,2}$ $\left[0.35,0.4\right]$
$h_{int,3}$ $\left[1.8,2\right]$
$h_{int,4}$ $\left[2.8,3.2\right]$
$h_{int,5}$ $\left[6.3,6.9\right]$
$c$ $\left[0.1,0.23\right]$
$h_{ext}$ $\left[0.85,0.95\right]$
Table 4: Maximum rewards of the searches with different number of states in the region given in Table 3. For the best six state spectrum the mean OPE coefficient of the last state is of the same order of magnitude as its standard deviation. Therefore we conclude that the five states are sufficient to complete the spectrum.
Number of states Max 1st state reward $\left|\frac{\sigma(C_{last})}{\mathrm{Mean}(C_{last})}\right|$
@@
Next, we look at how close are the found points to the exact minimal model spectra. The $7$-dimensional search region contains many minimal model spectra, so to visualize the search result in a systematic way, we plot the relative error using the procedure mentioned in Appendix D. Below, we briefly mention how it is done.
To visualize the minimal model spectra together with the search results in this seven dimensional space, for every found point we assign the closest (by Euclidean distance) minimal model spectra and calculate its relative error in every direction of the search space. Then we look at the distribution of these relative errors at the beginning and end of the search and plot their means and standard deviations. The search results are shown in Fig. 4.
In order to see how effective the CMA-ES search is, we also plot the relative errors of the top $10\%$ of the highest reward points at the initialization of the search. The error bars of the first four states (three in the case of three state search) are plotted with the means of corresponding relative errors and an associated error bar of length $2\sigma$. From left to right, the colors blue, orange, green and red labels the results of random sampling (or before running the algorithm), search in space with $2$ less dimension, that with $1$ less dimension, and that with exact dimension with the reference theories, and they are ordered from left to right in each plot. We see that the CMA-ES search significantly reduces the relative error.
Figure 4: Search result of different truncated space where only processes with 55 internal states are present.
IV Conclusion and outlook
In this work we proposed a simple objective function for searching CFT spectra satisfying the crossing equation, that is based on the statistics of solutions of OPE coefficients across many different $z,\bar{z}$ points. This approach allows us to study solutions to the bootstrap equations for non-unitary CFTs, where conventional convex optimization approaches are no longer applicable. We implemented this objective function as fully GPU parallelized function githubrepo:nucftb and demonstrated its applications in 2d CFT search for exact and truncated spectra using CMA-ES as an optimizer.
There are many immediate improvements one can implement. Firstly, we can ensure the suppression of the error term, $e$, in eq. (4) by constraining the contribution of states from large scaling dimensions. In particular, we can incorporate the implications of light-cone limit in $d>2$ Simmons-Duffin:2016wlq; Kusuki:2018wpa, where we know the existence of double-twist operators that were hidden in $e$ originally. Incorporating spin and considering $d>2$ will in general involve searches in large dimensions. This is an excellent arena to apply machine learning. In particular, since our objective is a structured (always made of conformal blocks) but non-differentiable function there might be more efficient strategies, compared to the generic optimization methods like CMA-ES, that utilize certain features of the objective function. Therefore, one could try to apply meta learning approaches shala2020cmaes; vishnu2019rnnopt; lange2023discovering; chen2022learning; 2016arXiv160601885L to see if a neural network could make use of some specific features of our objective functions to learn efficient optimization strategies.
Beyond non-unitary CFTs with negative scaling dimensions, central charge and squared OPE coefficienc ts, our approach is applicable for complex CFTs, which is related to dS/CFT correspondence Strominger:2001pn. Also, our idea should be applicable to cases where positivity cannot be applied such as five-point bootstrap Poland:2023vpn.
Acknowledgments
We would like to thank Po-Chung Chen, Li-Yuan Chiang, Miranda Cheng, Chong-Sun Chu, David Poland and Ning Su for enlightening discussions. Y-t H thanks Riken iTHEMS and the Yukawa Institute for Theoretical Physics at Kyoto University. Discussions during “Progress of Theoretical Bootstrap” were useful in completing this work. Y-tH is supported by the Taiwan National Science and Technology Council grant 112-2628-M-002-003-MY3 and 114-2923-M-002-011-MY5. S-C Lee is supported by 112-2628-M-002-003-MY3. H Liao is supported in part by the Ministry of Science and Technology grant 112-2112-M-002-024-MY3 and 112-2628-M-002-003-MY3.
Appendix A 2d Minimal Model CFT
@@ Acknowledgments
A special property of Virasoro minimal models is that their spectrum contains only finitely many Virasoro primary operators. Therefore, in principle using exact Virasoro conformal blocks we could solve (3) exactly. In practice, however, there is still error due to a numerical evaluation of the Virasoro blocks using finite recursion iterations.
Following Chen_2017 we calculate the Virasoro blocks
@@
Appendix B Selection of minimal models and comparison with search results
@@
In the following Table 5, we present examples of best rewards for four-state search in the region of table 3, where known five states models are available.
@@
Table 5: Nearest Minimal Model ($M$) vs. Found Spectrum
@@
Appendix C Specs of search
$x\sim\mathcal{N}(\mu,\sigma^{2}\mathcal{C}),$ (20)
To put CMA-ES on the GPU, we adopt the solver from Evotorch. In particular, we use the one in evotorch.algorithms.CMAES, which wraps the one in pycma so that it can be executed on the GPU backend. For the evaluation of the reward, we can parallelize the computation across $N_{c}$, $N_{z}$ and $N$ internal states. That is, during each epoch, we have $N_{c}$ candidates, each of which has $N_{z}\times N$ components in $G_{k,j}$ in (5), and we calculate these $N_{c}\times N_{z}\times N$ terms with a single pass through the GPU. Hence, we can significantly reduce the runtime.
Since CMA-ES is a self-adaptive method, it is almost parameter-free once we set the initial conditions (to start the search) and stopping condition (to prevent time waste). In fact, we only need to choose (1) an initial step size $\sigma_{0}$ reflecting the scale of the search space, (2) the population size $N_{p}$, (3) the maximum number of steps for evolution (also known as generation) $N_{g}$, and (4) the tolerance $t_{\sigma}$ to accept a solution when $\sigma$ is below this value. The initial mean $\mu_{0}$ is random sampled from the search space while the covariance matrix $\mathcal{C}$ is initialized to identity. The code and the setting we use to run the search are provided in githubrepo:nucftb.
In our case, we run the code githubrepo:nucftb on a single Nvidia A5500. As mentioned previously, for each epoch, the runtime is only a single pass on the GPU as long as there are enough threads for the computation. In our experience, GPU threads on A5500 are not exhausted when we calculate with up to $(N_{c},N_{z},N)=(500,400,5)$. In such case, with the level of recursion for Virasoro block evaluation being $40$, the runtime is about $2$ hours to search for $1000$ points. As long as we have sufficient number of GPU threads, the runtime does not scale with $N_{c}$, $N_{z}$ and $N$, and scale linearly with the recursion level and number of points to search.
Appendix D Procedure of error analysis

### v2

@@ Appendix A Supplemental tables
Table 2: Search window for the $3$-state Yang-Lee-region scan.
@@
Table 3: Search window for the five-state truncated search.
@@
$h_{int,1}$ $[0.23,0.3]$
$h_{int,2}$ $[0.35,0.4]$
$h_{int,3}$ $[1.8,2]$
$h_{int,4}$ $[2.8,3.2]$
$h_{int,5}$ $[6.3,6.9]$
$c$ $[0.1,0.23]$
$h_{ext}$ $[0.85,0.95]$
Table 4: Maximum rewards of the searches with different numbers of states in the region given in Table 3. For the best six-state spectrum, the mean OPE coefficient of the last state is of the same order as its standard deviation, indicating that five states are sufficient to complete the spectrum.
Number of states Max $N_{r}=1$ reward $\left|\frac{\sigma(C_{\rm last})}{\mathrm{Mean}(C_{\rm last})}\right|$
@@
Table 5: Initial search window for the non-unitary $c>1$ scan.
Variable Search bounds
$h_{int,1}$ [0.5,0.3][-0.5,-0.3]
$h_{int,2}$ $[0.1,0.3]$
$h_{int,3}$ $[0.3,0.5]$
$h_{int,4}$ $[0.5,0.7]$
$h_{int,5}$ $[0.7,0.9]$
$h_{int,6}$ $[0.9,1.1]$
$h_{int,7}$ $[1.1,1.3]$
$h_{int,8}$ $[1.3,1.5]$
$h_{int,9}$ $[1.5,1.7]$
$h_{int,10}$ $[1.7,1.9]$
$h_{int,11}$ $[1.9,2.1]$
$c$ $[1.05,1.25]$
$h_{ext}$ [0.25,0.05][-0.25,-0.05]
Table 6: Refined search window for the $12$- and $13$-state scans.
Variable Search bounds
$h_{int,1}$ [0.4,0.3][-0.4,-0.3]
$h_{int,2}$ $[0.25,0.35]$
$h_{int,3}$ $[0.4,0.5]$
$h_{int,4}$ $[0.55,0.65]$
$h_{int,5}$ $[0.75,0.85]$
$h_{int,6}$ $[0.95,1.05]$
$h_{int,7}$ $[1.15,1.25]$
$h_{int,8}$ $[1.3,1.4]$
$h_{int,9}$ $[1.55,1.65]$
$h_{int,10}$ $[1.75,1.85]$
$h_{int,11}$ $[2.0,2.1]$
$h_{int,12}$ $[2.1,3.5]$
$h_{int,13}$ $[2.9,4.5]$
$c$ $[1.01,1.11]$
$h_{ext}$ [0.17,0.07][-0.17,-0.07]
Table 7: Spectrum and inferred OPE coefficients for the highest-reward $k=13$ solution, with $(c,h_{ext})=(1.07662761,-0.11844774)$.
$j$ $h_{int,j}$ $C_{j}$
1 0.30347016-0.30347016 $1.02405595$
2 $0.33772710$ 0.11433376-0.11433376
3 $0.48003697$ $2.720435150$
4 $0.56931245$ 6.10600342-6.10600342
5 $0.82098216$ $12.38668966$
6 $1.03021669$ 21.7457633-21.7457633
7 $1.22721684$ $25.75052912$
8 $1.37006652$ 15.6260878-15.6260878
9 $1.60408115$ $3.273236841$
10 $1.82665527$ 0.6093234-0.6093234
11 $2.04846001$ $0.07390606$
12 $2.84190083$ 0.00008038-0.00008038
13 $3.21548438$ 0.0000190-0.0000190
Appendix B Specs of search
$x\sim\mathcal{N}(\mu,\sigma^{2}\mathcal{C}),$ (12)
To put CMA-ES on the GPU, we adopt the solver from Evotorch. In particular, we use evotorch.algorithms.CMAES, which wraps the one in pycma so that it can be executed on the GPU backend. For the evaluation of the reward, we can parallelize the computation across $N_{c}$, $N_{z}$ and $N$ internal states. That is, during each epoch, we have $N_{c}$ candidates, each of which has $N_{z}\times N$ components in $G_{\{z\}}$ in (5), and we calculate these $N_{c}\times N_{z}\times N$ terms with a single pass through the GPU. Hence, we can significantly reduce the runtime.
Since CMA-ES is a self-adaptive method, it is almost parameter-free once we set the initial conditions (to start the search) and stopping condition (to prevent time waste). In fact, we only need to choose (1) an initial step size $\sigma_{0}$ reflecting the scale of the search space, (2) the population size $N_{p}$, (3) the maximum number of steps for evolution (also known as generation) $N_{g}$, and (4) the tolerance $t_{\sigma}$ to accept a solution when $\sigma$ is below this value. The initial mean $\mu_{0}$ is random sampled from the search space while the covariance matrix $\mathcal{C}$ is initialized to identity. The code and the setting we use to run the search are provided in 1.
In our case, we run the code 1 on a single Nvidia A5500. As mentioned previously, for each epoch, the runtime is only a single pass on the GPU as long as there are enough threads for the computation. In our experience, GPU threads on A5500 are not exhausted when we calculate with up to $(N_{c},N_{z},N)=(500,400,5)$. In such case, with the level of recursion for Virasoro block evaluation being $40$, the runtime is about $2$ hours to search for $1000$ points. As long as we have sufficient number of GPU threads, the runtime does not scale with $N_{c}$, $N_{z}$ and $N$, and scale linearly with the recursion level and number of points to search.
Appendix C 2d Minimal Model CFT
@@ Appendix B Specs of search
A special property of Virasoro minimal models is that their spectrum contains only finitely many Virasoro primary operators. Therefore, in principle using exact Virasoro conformal blocks we could solve (LABEL:eq:_Crossing) exactly. In practice, however, there is still error due to a numerical evaluation of the Virasoro blocks using finite recursion iterations.
Following Chen et al. (2017) we calculate the Virasoro blocks
@@
Appendix D Further results
In this section, we collect additional results discussed in the main text. Firstly, we show projections of results for 2 and 3 state searches around Yang-Lee model. Each projection is compared with known minimal models, which are aligns well.
Here we present the top $5$ spectrum with $13$ states in terms of rewards. The resulting crossing violation is displayed in Fig 6.
Table 8: Examples of found 13 state spectra with $N_{r}=1$ reward greater than 10
$h_{\text{int},1}$ $h_{\text{int},2}$ $h_{\text{int},3}$ $h_{\text{int},4}$ $h_{\text{int},5}$ $h_{\text{int},6}$ $h_{\text{int},7}$ $h_{\text{int},8}$ $h_{\text{int},9}$ $h_{\text{int},10}$ $h_{\text{int},11}$ $h_{\text{int},12}$ $h_{\text{int},13}$ $c$ $h_{\text{ext}}$ $N_{r}=1$ reward
-0.303 0.338 0.480 0.569 0.821 1.030 1.227 1.370 1.604 1.827 2.048 2.842 3.215 1.077 -0.118 12.516
-0.301 0.315 0.476 0.593 0.792 1.007 1.210 1.358 1.614 1.802 2.061 2.872 3.621 1.071 -0.118 12.411
-0.300 0.349 0.472 0.563 0.760 0.992 1.213 1.375 1.616 1.792 2.069 2.855 3.881 1.078 -0.118 12.331
-0.303 0.344 0.471 0.608 0.778 1.028 1.177 1.364 1.614 1.789 2.019 2.879 3.410 1.079 -0.118 12.328
-0.306 0.343 0.469 0.585 0.785 0.972 1.208 1.357 1.622 1.802 2.094 2.833 4.093 1.078 -0.120 12.134
Appendix E Selection of minimal models and comparison with search results
@@
To illustrate the result of the selection process, we present five representative examples of identified Nearest Minimal model spectra, including the corresponding $(p,q,r,s)$ parameters, Found spectra, and the relative errors. The following Table 9 represents the results obtained by truncating a five internal states process to four internal states in the region of Table 3, where known five states models are available.
Table 9: Nearest Minimal Model ($M$) vs. Found Spectrum
@@
Appendix F Procedure of error analysis